# Time-Series Forecasting and Sequence Learning Using Memristor-based Reservoir System

### v1

@@ 5.3. Network Lifespan
The network lifespan (LSP) is defined by its sustainability for learning and acquiring new knowledge (Zyarah et al., 2020). In memristor-based architectures, the lifespan is highly affected by the memristor’s limited endurance, leading either to a gradual or severe drop in the dynamic range of network parameters, loss of elasticity, and eventually network performance. Typically, the memristor devices, particularly the oxide-based devices, offer an endurance ranging between $10^{6}-10^{12}$ (Coll et al., 2019). This low endurance is sufficient to get the network to continuously learn and update, but not for a long period of time.
Estimating the lifespan of the memristor-based network is not a trivial task as it is impacted by several factors, such as memristor endurance variability, input changing statistics, network convergence time, etc. However, in this work, first-order estimation for the lifespan will be considered as given in (13), where $E_{d}$ and $\sigma$, respectively, are the device endurance and variability. $U_{f}$ denotes the frequency update of the memristor devices while learning:
@@
For the used benchmarks, the PJM energy dataset, the input samples are recorded and presented to the network every hour. Thus, the lifespan of the proposed ESN accelerator when trained on the PJM energy dataset is $\sim$115,740 years, given $E_{d}=10^{9}$. However, this number can drop linearly with any increase in the update frequency. For instance, in Macky-Glass if the update occurs every 100ms, the lifespan of the same network drops to 3.21 years. Thus, we suggest the following techniques to enhance the lifespan: i) Use two memristors in differential configuration to model the synaptic weights. This not only results in a wide weight dynamic range, but it also helps extending the lifetime of the network, assuming the training process is conducted in an alternating manner. ii) Utilize R-M configuration to limit the voltage drop across memristor devices during the forming process and conventional switching. Reducing the voltage drop can either be achieved by explicitly integrating the memristor to a proper resistor (Kim et al., 2016) or inherently in 1T1M configuration. Integrating memristor devices into a crossbar structure is also expected to extend the device endurance due to the wire parasitic resistance (Salahuddin et al., 2018). Due to the simplicity and effectiveness of the first approach, alternating tuning, it is adopted in the developed ESN accelerator. Consequently, $\sim$2$\times$ enhancement in the lifespan of the network is achieved.
@@ 5.4. Latency
The latency, the time required to process each input sample presented to the memristor-based ESN accelerator, is calculated while processing the univariate time-series dataset. During forecasting, the latency (worst-case scenario) is estimated to be 45.83ns, unequally contributed by several units in the network, see Fig. 8-(a). However, the units that account $\sim$58of the delay are the neuron circuits which are implemented using Op-Amps to capture the non-linearity. This issue manifests only in the output layer as the reservoir neurons leverage the diode inherent characteristics to realize the non-linear activation function.
When comparing the presented memristor-based ESN accelerator with the digital counterpart clocked at 50MHz, 607$\sim$ 607$\times$ 607 × reduction in latency is witnessed when performing vector-matrix multiplication, owing to the extensive parallelism and in-memory computing of crossbar architecture, and 4$\sim$ 4$\times$ 4 × reduction in applying the non-linearity. It is important to mention that the latency of the units, developed under the 65nm process, is estimated in Cadence Virtuoso-ADE while considering resistance and capacitance parasitics, extracted in Mentor Graphics-Calibre.
@@ 5.5. Energy-delay product
The energy-delay product (EDP) of the proposed memristor-based accelerator implemented under 65nm process is 174.8nJ×$\times$×ns. It is estimated in Cadence Virtuoso while forecasting the load energy consumption from the PJM energy dataset. It is important to mention here that estimating the energy-delay product of a large-scale mixed-signal system tends to be challenging due to the disparity in time-scale between simulation and actual hardware models. For instance, the samples of the PJM dataset are recorded every hour and estimating the energy consumption using the same time-scale can take days or weeks. Thus, to speed up the process, the time is scaled down by $2.7\times 10^{-9}$. Fig. 8-(b) depicts the power consumption of the proposed ESN accelerator recorded over time. It can be noticed that the static power is dominating as compared to the dynamic power consumption which manifests when input samples are presented to the network. In Fig. 8-(c), we report the breakdown of the average power consumption among the main units. One may observe that most of the power budget is directed towards the Op-Amps. Thus, in this work, we strive to reduce the number of the used Op-Amp in the reservoir neurons by 60%. This results in 2$\times$ reduction in power consumption.
It is essential to highlight that the joint enhancements in power consumption, latency, and architecture are reflected on the entire system leading to 246$\times$ reduction in energy consumption of the proposed memristive ESN accelerator as compared to the digital counterpart (Zyarah et al., 2023). This makes the proposed accelerator more suitable to edge devices with stringent resources. Table 3 provides a high-level comparison of the proposed ESN accelerators with the previous works. Our accelerator offers in-situ training, allowing for: i) processing of stationary and non-stationary data, ii) hardware-aware learning and faster adaptation. Additionally, it features low latency and reasonable power consumption [^5] No comparison in terms of energy or energy-delay product is conducted due to the lack of information in the references above. as compared to other implementations reported in the literature.
@@ 6. Conclusions
In this paper, we propose an energy-efficient memristor-based ESN accelerator to enable time-series data processing and learning on edge devices with stringent resources. The proposed design features in-situ training, enabling processing of stationary and non-stationary data, hardware-aware learning, and faster adaptation. When evaluated for time-series forecasting using standard benchmarks, it is found that the hardware model experiences a marginal degradation in performance as compared to the software counterpart. This is attributed to several reasons, among which is the memristor devices’ non-idealities, limited precision of the used ADC, and undesired leakage in the sample and hold circuits. Regarding the lifespan of the network, we suggest alternating training approach to enhance the lifespan of the ESN. In case of latency while forecasting, we observe a significant reduction in latency as compared to the digital implementations. This is because the most extensive operations, multiply-accumulate, are performed concurrently and in-memory. When it comes to power and energy consumption, we notice that most of the power budget is directed towards the activation functions modeled by Op-Amps. Thus, we strive to simplify the implementation of the activation functions by using diode-based limiters. This results in 2$\times$ reduction in power consumption. The combined enhancements in architecture, latency, and power consumption give rise to a 246$\times$ reduction in energy consumption of the proposed memristive ESN accelerator as compared to the digital counterpart implemented at the same technology node.

### v2

@@ 5.3. Network Lifespan
The network lifespan (LSP) is defined by its ability to sustain learning and acquire new knowledge (Zyarah et al., 2020). In memristor-based architectures, the lifespan is highly affected by the memristor’s limited endurance, leading to either a gradual or severe drop in the dynamic range of network parameters, loss of elasticity, and eventually network performance. Typically, memristor devices, particularly oxide-based devices, offer an endurance ranging between $10^{6}-10^{12}$ (Coll et al., 2019). This low endurance is sufficient to get the network to continuously learn and update, but not for a long period of time.
Estimating the lifespan of the memristor-based network is not a trivial task as it is influenced by several factors, such as memristor endurance variability, changing input statistics, network convergence time, etc. However, in this work, first-order estimation for the lifespan will be considered as given in (13), where $E_{d}$ and $\sigma$, respectively, are the device endurance and variability. $U_{f}$ denotes the update frequency of the memristor devices while learning:
@@
For the used benchmarks, the PJM energy dataset, the input samples are recorded and presented to the network every hour. Thus, the lifespan of the proposed ESN accelerator, when trained on the PJM energy dataset, is $\sim$115,740 years, given $E_{d}=10^{9}$. However, this number can drop linearly with any increase in the update frequency. For instance, in Mackey-Glass if the update occurs every 100ms, the lifespan of the same network drops to 3.21 years. Thus, we suggest the following techniques to enhance the lifespan: i) Use two memristors in differential configuration to model the synaptic weights. This not only results in a wide weight dynamic range, but it also helps extending the lifetime of the network, assuming the training process is conducted in an alternating manner. ii) Utilize an R-M configuration to limit the voltage drop across memristor devices during the forming process and conventional switching. Reducing the voltage drop can either be achieved by explicitly integrating the memristor with a proper resistor (Kim et al., 2016) or inherently in a 1T1M configuration. Integrating memristor devices into a crossbar structure is also expected to extend the device endurance due to the wire parasitic resistance (Salahuddin et al., 2018). Due to the simplicity and effectiveness of the first approach, alternating tuning, it is adopted in the developed ESN accelerator. Consequently, $\sim$2$\times$ enhancement in the lifespan of the network is achieved.
@@ 5.4. Latency
The latency, the time required to process each input sample presented to the memristor-based ESN accelerator, is calculated while processing the univariate time-series dataset. During forecasting, the latency (worst-case scenario) is estimated to be 45.83ns, unequally contributed by several units in the network, see Fig. 6-(a). However, the units that account $\sim$58of the delay are the neuron circuits, which are implemented using two Op-Amps to capture the non-linearity. This issue manifests only in the output layer, as each reservoir neuron uses one Op-Amp and a leakage cell to capture the non-linearity and leaky-integrated feature, respectively.
When comparing the presented memristor-based ESN accelerator with the digital counterpart clocked at 50MHz, $\sim$607$\times$ reduction in latency is witnessed when performing vector-matrix multiplication, owing to the extensive parallelism and in-memory computing of crossbar architecture, and $\sim$4$\times$ reduction in applying the non-linearity. It is important to mention here that to ensure precision when estimating the latency of the developed ESN accelerator, we built a large-scale network (1$\times$105$\times$1) in Cadence Virtuoso and estimated the time required to propagate an input signal through the individual components (developed under the 65nm process) and layers. The estimation of the propagation time is done using Cadence Virtuoso-ADE while considering resistance and capacitance parasitics, extracted in Mentor Graphics-Calibre.
@@ 5.5. Energy-delay product
The energy-delay product (EDP) of the proposed memristor-based accelerator implemented using a 65nm process is 153.68nJ×$\times$×ns. It is estimated in Cadence Virtuoso while forecasting the load energy consumption from the PJM energy dataset. It is important to note here that estimating the energy-delay product of a large-scale mixed-signal system tends to be challenging due to the disparity in the time-scale between simulation and actual hardware models. For instance, the samples of the PJM dataset are recorded every hour and estimating the energy consumption using the same time-scale can take days or weeks. Thus, to speed up the process, the time was scaled down by $2.7\times 10^{-9}$. Fig. 6-(b) depicts the power consumption of the proposed ESN accelerator recorded over time. It can be observed that the static power is dominating as compared to the dynamic power consumption which manifests when input samples are presented to the network. In Fig. 6-(c), we report the breakdown of the average power consumption among the main units. One may observe that most of the power budget is directed towards the Op-Amps. Thus, in this work, we strive to reduce the number of the Op-Amp used in the reservoir layer, which results in more than 2$\times$ reduction in power consumption.
It is essential to highlight that the joint enhancements in power consumption, latency, and architecture are reflected on the entire system leading to 247$\times$ reduction in energy consumption of the proposed memristive ESN accelerator compared to the digital counterpart (Zyarah et al., 2023). This makes the proposed accelerator more suitable for edge devices with stringent resources. Table 3 provides a high-level comparison of the proposed ESN accelerator with the previous works. Our accelerator offers in-situ training, allowing for: i) processing of stationary and non-stationary data locally on the device, ii) hardware-aware learning and faster adaptation. Additionally, it features low latency as well as reasonable power consumption [^8] No comparison in terms of energy or energy-delay product is conducted due to the lack of information in the references above. as compared to other implementations reported in the literature.
@@ 6. Conclusions
In this paper, we propose an energy-efficient memristor-based ESN accelerator to enable time-series data processing and learning on edge devices with stringent resources. The proposed design features in-situ training, enabling processing of stationary and non-stationary data, hardware-aware learning, and faster adaptation. When evaluated for time-series forecasting using standard benchmarks, it was found that the hardware model experiences a marginal degradation in performance compared to the software counterpart. This is attributed to several reasons, among which are the memristor devices’ non-idealities, limited precision of the used ADC, and undesired leakage in the sample and hold circuits. Regarding the lifespan of the network, we suggest an alternating training approach to enhance the lifespan of the ESN. In the case of latency, while forecasting, we observe a significant reduction in latency compared to the digital implementations. This is because the most extensive operations, multiply-accumulate, are performed concurrently and in memory. When it comes to power and energy consumption, we notice that most of the power budget is directed toward the Op-Amps used to model the activation functions and capture the leaky-integrated feature. Thus, we strive to reduce the number of Op-Amps utilized to emulate the activation functions and capture the leaky-integrated features via introducing the memristor-based leakage cell. This results in more than 2$\times$ reduction in power consumption. The combined enhancements in architecture, latency, and power consumption give rise to a 247$\times$ reduction in energy consumption of the proposed memristive ESN accelerator compared to the digital counterpart implemented at the same technology node.